# Unleashing the Power of LLMs as Multi-Modal Encoders for Text and Graph-Structured Data

### v1

@@ C.2.1. KG-Contextualized QA
(8) $\mathcal{L}^{\rm(task)}=-\sum_{i=1}^{n_{\rm batch}}\sum_{j=1}^{n_{\rm choice}}y_{j}^{(i)}\log(\hat{y}_{j}^{(i)})$
@@
C.2.2. Graph-Text Pair Classification Tasks
@@
C.2.3. Retrieval Tasks
For retrieval tasks, we apply infoNCE loss (Oord et al., 2018) for training. Specifically, given a batch of positive pairs $(d_{i},p_{i})$, we assume that $(d_{i},p_{i})$ is a positive pair and $(d_{i},p_{j})$ for $i\neq j$ a negative pair. By applying infoNCE loss, we have
(10) $\mathcal{L}^{(task)}=-\sum_{i=1}^{n_{\text{batch}}}\log\left(\frac{e^{\text{sim}(a_{i},p_{i})/\tau}}{e^{\text{sim}(a_{i},p_{i})/\tau}+\sum_{j\neq i}e^{\text{sim}(a_{i},p_{j})/\tau}}\right)$
@@ C.2.3. Retrieval Tasks
Table 8. Ablation Study on different branches of Janus: This table reports the test accuracy of Janus when using only the graph + text branch or only the graph description + text branch.
@@
Janus 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10
Table 9. The effect of graph encoder choice on test accuracy.
@@
GCN (Kipf and Welling, 2017) 80.95 $\pm$ 0.28 84.67 $\pm$ 1.27
GAT (Velickovic et al., 2018) 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10
GIN (Xu et al., 2019) 80.23 $\pm$ 0.57 85.40 $\pm$ 0.92
Appendix D Experiment Details
Table 10. Hyperparameters for training Janus.
@@ Appendix D Experiment Details
D.1. Additional Implementation Details
We implement models using PyTorch (Paszke et al., 2019) and PyTorch Geometric (Fey and Lenssen, 2019). All experiments are conducted on a single A100 80GB GPU. For baselines that we implemented ourselves, we followed the settings and hyperparameters described in the original papers to ensure fair comparisons. Detailed hyperparameter settings for Janus can be found in Table 10. Additionally, since we freeze the parameters of the LLM backbone in Janus, the graph descriptions are encoded in advance before training. This strategy improves training efficiency and significantly reduces memory usage. For the E5-Mistral (Wang et al., 2023) baseline, we freeze the parameters and only train the classification head, while for other LMs we update all parameters.
D.2. Additional Results
D.2.1. Effect of Different Branches in Janus
To better understand the contributions of each branch in Janus, we perform an ablation study, as shown in Table 8. This analysis evaluates the performance of Janus when using only the graph + text branch or only the graph description + text branch and compares it to the full version of Janus, which integrates both branches. We find that while each branch individually yields comparable performance, the combination of both branches—enhanced by our graph-text alignment technique—leads to significant performance gains. This demonstrates the effectiveness of our proposed methods.
D.2.2. Effect of Graph Encoder Choice
We also investigate the impact of different graph encoders on the performance of Janus, as shown in Table 9. We observe that the performance with GAT achieves the best results on both CommonsenseQA and OpenBookQA.

### v2

@@ D.2.1 KG-Contextualized QA
$\mathcal{L}^{\rm(task)}=-\sum_{i=1}^{n_{\rm batch}}\sum_{j=1}^{n_{\rm choice}}y_{j}^{(i)}\log(\hat{y}_{j}^{(i)})$ (7)
@@
D.2.2 Graph-Text Pair Classification Tasks
@@
D.2.3 Retrieval Tasks
For retrieval tasks, we apply infoNCE loss Oord et al. (2018) for training. Specifically, given a batch of positive pairs $(d_{i},p_{i})$, we assume that $(d_{i},p_{i})$ is a positive pair and $(d_{i},p_{j})$ for $i\neq j$ a negative pair. By applying infoNCE loss, we have
@@ D.2.3 Retrieval Tasks
Table 5: Ablation Study on different branches of GT2Vec: This table reports the test accuracy of GT2Vec when using only the graph + text branch or only the graph description + text branch.
@@
GT2Vec 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10
Table 6: The effect of graph encoder choice on test accuracy.
@@
GCN Kipf and Welling (2017) 80.95 $\pm$ 0.28 84.67 $\pm$ 1.27
GAT Velickovic et al. (2018) 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10
GIN Xu et al. (2019) 80.23 $\pm$ 0.57 85.40 $\pm$ 0.92
Table 7: LoRA Hyperparameters
Hyperparameter Value
Rank ($r$) 8
Alpha ($\alpha$) 16
Dropout 0.05
Target Modules {q_proj, k_proj, v_proj, o_proj,
gate_proj, up_proj, down_proj}
Appendix E Experiment Details
Table 8: Hyperparameters for training GT2Vec.
@@ Appendix E Experiment Details
E.1 Additional Implementation Details
We implement models using PyTorch Paszke et al. (2019) and PyTorch Geometric Fey and Lenssen (2019). All experiments are conducted on a single A100 80GB GPU. For baselines that we implemented ourselves, we followed the settings and hyperparameters described in the original papers to ensure fair comparisons. Detailed hyperparameter settings for GT2Vec can be found in Table 7 and Table 8.
E.2 Additional Results
E.2.1 Effect of graph context information.
When GT2Vec is input with only textual input, excluding graph context information, there is a significant drop in performance, with accuracy decreasing to 69.49% on CommonsenseQA and 74.80% on OpenBookQA. This sharp decline underscores the critical role that graph context plays in enhancing the model’s ability to understand and reason with the data. The graph provides structured knowledge that complements the unstructured text, and without it, the model relies solely on the textual input, which limits its capacity to effectively handle complex embedding tasks.
E.2.2 Effect of Graph-Text Alignment.
Removing contrastive learning from GT2Vec leads to a performance drop, with the accuracy decreasing from 81.09% to 79.69% on CommonsenseQA, and from 86.67% to 84.80% on OpenBookQA. This highlights the importance of contrastive learning in aligning graph and text embeddings, ensuring better integration of multimodal information.
Table 9: Ablation study results on test accuracy.
Methods CommonsenseQA OpenBookQA
GreaseLM Zhang et al. (2022) 74.20 $\pm$ 0.40 83.87 $\pm$ 1.29
GT2Vec w/o graph 69.49 $\pm$ 0.28 74.80 $\pm$ 0.35
GT2Vec w/o alignment 79.69 $\pm$ 0.85 84.80 $\pm$ 1.06
GT2Vec 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10
E.2.3 Effect of Different Branches in GT2Vec
To better understand the contributions of each branch in GT2Vec, we perform an ablation study, as shown in Table 5. This analysis evaluates the performance of GT2Vec when using only the graph + text branch or only the graph description + text branch and compares it to the full version of GT2Vec, which integrates both branches. We find that while each branch individually yields comparable performance, the combination of both branches—enhanced by our graph-text alignment technique—leads to significant performance gains. This demonstrates the effectiveness of our proposed methods.
E.2.4 Effect of Graph Encoder Choice
We also investigate the impact of different graph encoders on the performance of GT2Vec, as shown in Table 6. We observe that the performance with GAT achieves the best results on both CommonsenseQA and OpenBookQA.
E.2.5 Impact of Node Embedding Initialization
Table 10: Impact of different graph node embedding initialization models on accuracy for the CommonsenseQA dataset.
Model Output Dimension Accuracy
RoBERTa-Large Liu et al. (2019) 1,024 81.09 $\pm$ 0.73
E5-Small Wang et al. (2022) 384 80.37 $\pm$ 0.38
E5-Base Wang et al. (2022) 768 80.77 $\pm$ 0.05
E5-Large Wang et al. (2022) 1,024 80.79 $\pm$ 0.30
E5-Mistral Wang et al. (2023) 4,096 80.42 $\pm$ 0.90
Interestingly, even when using lighter models for node embedding initialization, such as E5-Small, which has a smaller output dimension (384), the performance remains competitive. This demonstrates that even when the node initialization model differs from the actual backbone used in GT2Vec, there is no significant performance degradation. For example, compared with E5-Mistral node embedding initialization, the lighter E5-Small model offers the advantage of reduced computational cost without compromising much on accuracy.